OpenStudy (anonymous):

Equations? Thanks!!! :) 1. Which ratio is equal to 5/6? 2. Which ratio is not equal to 4/9 3. Solve the proportion. 16/17 = x/40 4. Solve for a. 4a/9 = 100/45 5. Solve for b. 3/0.8 = 12/b 6. Solve. 3/4 = 21/2+d 7. When surveyed five out of every eight students liked food in the cafeteria. If 60 liked the food how many were surveyed?

OpenStudy (anonymous):

For 1. and 2. what are the options to choose from?? 3. Solve the proportion. 16/17 = x/40 (16 x 40)/17 = x so, x = 640/17 4. Solve for a. 4a/9 = 100/45 a = (100 x 9)/(45 x 4) a = 900/180 a = 5 You can solve 5. and 6. by method used for 3. and 4. 5. Solve for b. 3/0.8 = 12/b 6. Solve. 3/4 = 21/2+d 7. When surveyed five out of every eight students liked food in the cafeteria. If 60 liked the food how many were surveyed? 5/8 :: 60/x 5x = 60 x 8 x = 480/5 x = 96 so 96 students were surveyed